The equation of an estimated demand function is as follows: -

QdA (Quantity demand for A) = 200.5 - 2.5 Pa - 1.5Pb + 3.5 I

where, Pa = Price of A

Pb = Price of B [It is a related product]

I = Income

(i) Determine the demand when Pa = $ 500, Pb = $ 100 and I = $ 3000

(ii) Estimate the price elasticity, cross-price elasticity and income elasticity of the demand according to point method.

(iii) Estimate the elasticity of the demand according to proportion method if Pa2 = $ 515, Pb2 = $ 105 and I = $ 3500.


1
Expert's answer
2021-08-13T11:44:17-0400

Given

"Q^d_A= 200.5\u22122.5P_A \u22121.5P_B+3.5 I ...... (1)"

Where Price of A = Pa and Pb is the price of B and I is income 

 When Pa = $ 500, Pb = $ 100 and I = $ 3000

Then 

"Q^d_A=200.5−2.5(500)−1.5(100)+3.5(3000)\\⇒ Q^d_A=9300.5"

(I)

Price elasticity of the demand 

Differentiate equation 1 w r t PA

"\frac{dQ^d_A}{dP_A}= −2.5"

Price elasticity(E)"=\frac{dQ^d_A}{dP_A}×\frac{P_A}{Q^d }"

"⇒E=−2.5×\frac{500}{9300.5}\\⇒E=−0.1344\\⇒|E|=0.1344"


(ii)

Cross elasticity

Differentiate equation 1 w r t PB

"\frac{dQ^d_A}{dP_B}= −1.5"

Cross-Price elasticity"=\frac{dQ^d_A}{dP_B}×\frac{P_B}{Q^d }"

"EP_B=−1.5×\frac{100}{9300.5}\\⇒EP_B=−0.0161"


(iii)

Income elasticity:

Differentiate equation 1 w r t I

"\frac{dQ^d_A}{dI}= 3.5"

"Income \space elasticity = \frac{dQ^d_A}{dI}×\frac{IQ}{A^d}"

"= 3.5×\frac{3000}{9300.5}\\=1.129"